OpenStudy (anonymous):

solve 1/3 - 7/2t=-5/6

OpenStudy (anonymous):

t = 3

OpenStudy (anonymous):

how do u work that out

OpenStudy (anonymous):

Get a common denominator if 6t.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

2t/6t - 21/6t = -5t/6t (2t-21)/6t = -5t/6t 2t-21 = -5t -21 = -7t t = 3